OpenStudy (anonymous):

What is the value of x in the equation 3(2x + 8) = 0? –8 –4 4 8

OpenStudy (anonymous):

-4 ?

OpenStudy (anonymous):

are u sure ?

OpenStudy (anonymous):

3 (2x+8)=0 First Distribute

OpenStudy (cgbork):

Answer should be 4

OpenStudy (anonymous):

The 3 with 2x and 8

OpenStudy (anonymous):

but u just said the asnwer should be negative 4 or at least thats what i got

OpenStudy (cgbork):

You should distribute. Niaa was right, I was wrong.

OpenStudy (cgbork):

3(2x+8)=0 6x+24=0 6x=-24 6x/6=24/6 x=4 Do you see it?

OpenStudy (anonymous):

Yea The Answer is -4 , yes yes @cgbork that's correct

OpenStudy (anonymous):

im so confused though

OpenStudy (cgbork):

sorry, x =-4. I keep forgetting the negative sign.

OpenStudy (anonymous):

ok thank you lol

OpenStudy (cgbork):

My apologies. If only you could edit your posts on here

OpenStudy (anonymous):

yea true its ok

OpenStudy (anonymous):

Ok First you Have To Distribute. When ever you have a number in front of a parentheses, you always distribute before solving. So 3 (2x+8)=0 You Distribute by MultiplYing 3 by 2x and 8 Which will give you 6x+24=0 Then You Subtract 24 from both sides of the equal sign Which will give you 6x=-24 Then You Just Divide By Putting a 6 Under Everything and So The 6 cancel out with its self and -24÷6 is -4 So you Will Be left with, which is your answer ×=-4